OpenStudy (anonymous):

Let f(x) = 5x – 26 and g(x) = 6x. Part 1: What is (f ◦ g)(x) ? Part 2: Use complete sentences to describe the method you used to solve this problem.

jimthompson5910 (jim_thompson5910):

(f ◦ g)(x) means f(g(x))

jimthompson5910 (jim_thompson5910):

f(x) = 5x – 26 f(g(x)) = 5(g(x)) – 26 ... replace all copies of x with g(x) f(g(x)) = 5(6x) – 26 ... replace all copies of g(x) on the right side with 6x f(g(x)) = 30x – 26

jimthompson5910 (jim_thompson5910):

and this is all substitution we're using

OpenStudy (anonymous):

so the answer is f(g(x)) = 30x – 26

jimthompson5910 (jim_thompson5910):

to part 1, yes

jimthompson5910 (jim_thompson5910):

part 2 is just explaining how you got part 1

OpenStudy (anonymous):

okay and i already know how to do(: thanks jim!

jimthompson5910 (jim_thompson5910):

sure thing